# The Curious Case of In-Training Compression of State Space Models

### v3

@@ 2.1 Discrete Linear Time Invariant Systems
The pair $({\bm{A}},{\bm{B}})$ is controllable, i.e., the state $\bm{h}$} can be be steered from any initial state to any final state in finite time.
@@ 2.1.1 Controllability and Observability Gramians
Under assumptions 2.1 and 2.2, there exist a unique symmetric positive definite solution ${\bm{P}}\in\mathbb{R}^{n\times n}$ to the discrete Lyapunov equation:
${\bm{A}}{\bm{P}}{\bm{A}}^{\text{T}}-{\bm{P}}+{\bm{B}}{\bm{B}}^{\text{T}}$ $=0,$ $\quad{\bm{P}}$ $=\sum_{i=0}^{\infty}{\bm{A}}^{i}{\bm{B}}{\bm{B}}^{\text{T}}({\bm{A}}^{\text{T}})^{i}$ (2)
@@
Under assumptions 2.1 and 2.3, there exist a unique symmetric positive definite solution ${\bm{Q}}\in\mathbb{R}^{n\times n}$ to the discrete Lyapunov equation:
${\bm{A}}^{\text{T}}{\bm{Q}}{\bm{A}}-{\bm{Q}}+{\bm{C}}^{\text{T}}{\bm{C}}$ $=0,$ $\quad{\bm{Q}}$ $=\sum_{i=0}^{\infty}({\bm{A}}^{\text{T}})^{i}{\bm{C}}^{\text{T}}{\bm{C}}{\bm{A}}^{i}$ (3)
@@ 2.1.2 Balanced Realizations
Any stable, minimal discrete LTI system admits a balanced realization, in which the controllability and observability Gramians coincide as ${\bm{W}}=\text{diag}({\bm{\sigma}})=\text{diag}(\sigma_{1},\dots,\sigma_{n})$, with $\sigma_{1}\geq\cdots\geq\sigma_{n}>0$ called “Hankel singular values“ (HSV). This diagonal balanced realization can be explicitly constructed .
@@ 2.2 Spectral stability of Hermitian matrices
Let $\bm{W}$} and $\bm{W}$$^{\prime}$ be Hermitian matrices of size $n$ (i.e. symmetric for real-valued matrices), and let $\delta{\bm{W}}={\bm{W}}^{\prime}-{\bm{W}}$. Also, $\forall i\in[1,\cdots,n]$, let $\sigma_{i}({\bm{W}})$ represent the $i$-th largest eigenvalue of $\bm{W}$}. The ordering $\sigma_{1}({\bm{W}})\geq\cdots\geq\sigma_{n}({\bm{W}})$ can always be established as the eigenvalues of Hermitian matrices are guaranteed to be real.
@@
$\forall i\in[1,\cdots,n]$, $\sigma_{i}({\bm{W}})$ is Lipschitz-continuous on the space of Hermitian matrices with operator norm:
$|\sigma_{i}({\bm{W}}^{\prime})-\sigma_{i}({\bm{W}})|\leq\max_{i=1,\cdots,n}(|\sigma_{i}(\delta{\bm{W}})|)=\max(|\sigma_{1}(\delta{\bm{W}})|,|\sigma_{n}(\delta{\bm{W}})|).$ (7)
@@ 3 The Proposed In-Training Reduction Scheme
Our general pipeline (illustrated in Figure 1) is designed to be applicable to all types of SSMs as it surgically acts on the dynamical systems within SSM layers of models regardless of the choice of projections, non-linear activations, convolutions, skip connection, etc.
@@ 3.1 CompreSSM: the algorithm
At an given training step, we proceed per block. For an input feature vector sequence ${\bm{x}}\in\mathbb{R}^{H\times L}$, where $H$ is the inner dimension and $L$ the sequence length, common SSM blocks contain either a single Multi-Input Multi-Output (MIMO) system that transforms the sequence ${\bm{x}}\in\mathbb{R}^{H\times L}$ to ${\bm{y}}\in\mathbb{R}^{H\times L}$, or $H$ independent per-channel Single-Input Single-Output (SISO) systems that transform ${\bm{x}}^{i}\in\mathbb{R}^{1\times L}$ to ${\bm{y}}^{i}\in\mathbb{R}^{1\times L}$, for $i\in{1,\cdots,H}$. In the latter case, we proceed per channel.
@@
    Find the smallest rank $r$ such that the top-$r$ singular values account for a predetermined threshold $\tau\in[0,1]$ of the total energy,
    $r=\min\Bigl\{k\in\{1,\cdots,n\}:\sum_{i=1}^{k}\sigma_{i}\;\geq\;\tau\sum_{i=1}^{n}\sigma_{i}\Bigr\},$ (8)
@@ 3.3 In-Training Reduction
Also recall that the Hankel singular values can be obtained as the square root of the eigenvalues of ${\bm{P}}{\bm{Q}}$ (Equation 4). Generally, ${\bm{P}}{\bm{Q}}$ need not be symmetric, but we use this form for computational efficiency. Noticing that ${\bm{P}}{\bm{Q}}$ is similar to the symmetric positive definite matrix ${\bm{P}}^{1/2}{\bm{Q}}{\bm{P}}^{1/2}$, we let ${\bm{H}}=\sqrt{{\bm{P}}^{1/2}{\bm{Q}}{\bm{P}}^{1/2}}$. The matrix $\bm{H}$}’s eigenvalues are exactly the Hankel singular values. In addition, by composition, $\bm{H}$} is continuous with respect to the perturbations to the system, so we write ${\bm{H}}^{\prime}={\bm{H}}+\delta{\bm{H}}$ with $\delta{\bm{H}}$ a continuous function of $(\delta{\bm{A}},\delta{\bm{B}},\delta{\bm{C}})$.
@@ 4.1 Experimental Setup
Reduction starts from the full model order reported in column 4 of Table 3. For all datasets but IMDB and sMNIST, we attempt four equidistant reduction steps during the warm-up period of the learning rate, which equals $10\%$ of the total steps. Doing so ensures maximal speed up potential for the subsequent $90\%$ of training, while also staying robust to large early training .
@@ 4.2 Empirical Results
However, on datasets where state dimension does correlate with model performance, CompreSSM improves small model performance. On CIFAR10, for example, model performance almost stays constant as a function of reduction tolerance (and thus for different state dimensions), while the non-compressed counterparts exhibit a approximately $10\%$ performance drop (see entry 1 in Table 1 or Figure 3(a)). The sMNIST results paint a similar picture (entry 5 in Table 1 or Figure 4(a)).
@@
On Pathfinder, we can observe a similar trend (Figure 4(e)), where the unreduced model performance does not show a strong correlation with the state dimension. Just for the smallest state dimension, the unreduced model performs sees a small drop in performance and is outperformed by the top-1 reduced model.
The IMDB results reveal the importance of the prerequisites mentioned in Section 3.1 (see Figure 4(d)). Even after significantly reducing the parameters and increasing the droprate (see Appendix A), the unreduced only learn for roughly $8k$ steps before overfitting. However, in order to do in-training balanced truncation, the actual training phase needs to be long enough to allow for a couple of reduction steps with a large enough spacing so that the model can follow the training dynamics for a while without being pruned. Indeed, for non-aggressive pruning (that is, pruning with small tolerance $\tau$), the top reduced models often outperform the baseline.
@@ 5 Related Work
Although not a compression approach, the authors in Yu et al. (2024) approach the problem of Hankel singular value efficiency from the perspective of SSM initialization, providing an approach to avoid fast decaying singular values.
@@ 6 Conclusion
In conclusion, we proposed a novel framework for principled compression of SSMs during training called CompreSSM. Drawing on classical control theory, we base our approach on balanced truncation. In particular, we show that Hankel singular values (HSVs), the corner stone of balanced truncation, preserve the rank of their dominant dimensions during training, which allows for safely truncating dimensions associated with smaller HSVs.
We test CompreSSM by training LRUs on a range of tasks of varying complexity. Provided there is a correlation between state dimension and model performance as well as enough training steps in between two reduction steps, we verify empirically that compressed models outperform their uncompressed counterparts while delivering better performance per unit of training time.
Finally, we provide an outlook on how to do in-training balanced truncation for linear time-varying systems. In this selective case, we propose averaging the dynamics over the input space and applying the reduction scheme to this time-independent system. For future work, we would like to extend CompreSSM to linear self-attention models, e.g., Gated Linear Attention (Yang et al., 2023), Mamba2 (Dao and Gu, 2024), and Gated DeltaNet (Yang et al., 2024), which is another class of models based on linear time-varying systems.
@@ A.1 Solving The Lyapunov Equations for Diagonal SSMs
${\bm{P}}_{ij}$ $=\frac{\left({\bm{B}}{\bm{B}}^{\top}\right)_{ij}}{1-\lambda_{i}\lambda_{j}},\quad{\bm{Q}}_{ij}=\frac{\left({\bm{C}}^{\top}{\bm{C}}\right)_{ij}}{1-\lambda_{i}\lambda_{j}}\quad\forall 1\leq i,j\leq n$ (14)
@@ C.1 Balanced Truncation Sanity Check
Figure 5 shows the average model accuracy on the validation dataset during training, and Figure 6 depicts the HSVs evolution for a single seed across the three reduction schemes. Removing the top HSVs clearly incurs catastrophic performance damage, that is not recoverable even with 90% of subsequent training. One can see th These drops in performance are present yet less pronounced when removing random HSVs. In this case, the original gains gathered from training at larger state dimension are not completely lost, as the network is somewhat able to recover to perform slightly better than the baseline trained with dimension 32 from the start. In the case of proper bottom HSV truncation, there are no notable drops in performance in comparison to the 256 baseline, apart from a small departure at the final reduction step. The performance observations are also backed up by the HSVs evolution in each case. Indeed, we observe that the correct balanced truncation maintains a majority concentration of large HSVs throughout training, all within a single order of magnitude, maximizing the per dimension contribution. Random truncation leads to a sparser spread of HSVs with many small surviving HSVs trailing an order of magnitude or two behind the leading dimensions. Expressivity is most diluted in the adversarial case, where removing the top values leads to dynamics with only a small proportion of HSVs carrying training while the largest bulk remains several orders of magnitude behind.
@@ D.2 Reduction Pipeline Complexity
For $\mathbf{B}\in\mathbb{R}^{n\times p}$, computing $\mathbf{B}\mathbf{B}^{\top}$ takes $\mathcal{O}(n^{2}p)$ FLOPS. Similarly, computing $\mathbf{C}^{\top}\mathbf{C}$ for $\mathbf{C}\in\mathbb{R}^{q\times n}$ takes $\mathcal{O}(n^{2}q)$ FLOPS. Thus, computing the Gramians $\bm{P}$} and $\bm{Q}$} requires $\mathcal{O}(n^{2}p+3n^{2})$ and $\mathcal{O}(n^{2}q+3n^{2})$ FLOPS, respectively.
@@
Step 4: finding the smallest rank $r$ that maintains at least $\tau$ of the total energy requires computing the cumulative sum of the HSV, which is $\mathcal{O}(n)$ FLOPS.
@@ E.3.2 Practical reduction workflow
From each surrogate we form controllability and observability Gramians ${\bm{P}}^{i}$ and ${\bm{Q}}^{i}$. Their product yields the Hankel singular values ${\bm{g}}^{i}$, which determine the retained dimensionality ${\bm{r}}^{i}$. We compute the balancing transform ${\bm{T}}^{i}$ and its inverse $({\bm{T}}^{i})^{-1}$ that simultaneously diagonalize the Gramians ${\bm{P}}^{i}$ and ${\bm{Q}}^{i}$.
Applying $({\bm{T}}^{i})^{-1}\widetilde{{\bm{A}}}^{i}{\bm{T}}^{i}$ and truncating to the first ${\bm{r}}^{i}$ rows/columns produces ${\bm{A}}_{\text{red}}^{i}$. Similarly,
${\bm{B}}_{\text{red}}^{i}=({\bm{T}}^{i})^{-1}\widetilde{{\bm{B}}}^{i}\big[:{\bm{r}}^{i}\big],\qquad{\bm{C}}_{\text{red}}^{i}=\widetilde{{\bm{C}}}\,{\bm{T}}^{i}\big[:,:{\bm{r}}^{i}\big].$
@@
We then keep only the first ${\bm{r}}^{i}$ coordinates before feeding the selective recurrence. The vector of ranks ${\bm{r}}=({\bm{r}}^{1},\dots,{\bm{r}}^{d_{\text{inner}}})$ is passed to the selective-scan CUDA kernels, which terminate their inner loops at ${\bm{r}}^{i}$ for channel $i$. This eliminates all work on pruned coordinates and makes the runtime cost proportional to the reduced dimension.
@@ E.3.3 Experimental Evaluation:
We test the performance of CompreSSM in this setting by compressing with $\tau=0.001$. This means the model retains at least $99.99\%$ of its energy at every reduction step. The final model has an average state dimension of $\sim 12$, starting from $128$. This shows that the HSV spectrum is very tail-heavy, with most energy in just a hand full of dimensions. This explains why random dropping of HSV (the red markers in Figure 18) performs competitively with balanced truncation (the orange markers), though at a higher variance.
Even though the correlation between state dimension and performance is not very strong, CompreSSM yields competitive performance on Mamba, at a lower variance. Especially when accounting for outliers. However, we believe the approach for LTI systems should be refined in later works to account for heavy-tailed distributions of HSV.

### v4

@@ 2.1 Discrete Linear Time Invariant Systems
The pair $({\bm{A}},{\bm{B}})$ is controllable, i.e., the state $\bm{h}$} can be steered from any initial state to any final state in finite time.
@@ 2.1.1 Controllability and Observability Gramians
Under assumptions 2.1 and 2.2, there exists a unique symmetric positive definite solution ${\bm{P}}\in\mathbb{R}^{n\times n}$ to the discrete Lyapunov equation:
${\bm{A}}{\bm{P}}{\bm{A}}^{\top}-{\bm{P}}+{\bm{B}}{\bm{B}}^{\top}$ $=0,$ $\quad{\bm{P}}$ $=\sum_{i=0}^{\infty}{\bm{A}}^{i}{\bm{B}}{\bm{B}}^{\top}({\bm{A}}^{\top})^{i}\,.$ (2)
@@
Under assumptions 2.1 and 2.3, there exists a unique symmetric positive definite solution ${\bm{Q}}\in\mathbb{R}^{n\times n}$ to the discrete Lyapunov equation:
${\bm{A}}^{\top}{\bm{Q}}{\bm{A}}-{\bm{Q}}+{\bm{C}}^{\top}{\bm{C}}$ $=0,$ $\quad{\bm{Q}}$ $=\sum_{i=0}^{\infty}({\bm{A}}^{\top})^{i}{\bm{C}}^{\top}{\bm{C}}{\bm{A}}^{i}\,.$ (3)
@@ 2.1.2 Balanced Realizations
Any stable, minimal discrete LTI system admits a balanced realization, in which the controllability and observability Gramians coincide as ${\bm{W}}=\text{diag}({\bm{\sigma}})=\text{diag}(\sigma_{1},\dots,\sigma_{n})$, with $\sigma_{1}\geq\cdots\geq\sigma_{n}>0$ called “Hankel singular values” (HSV). This diagonal balanced realization can be explicitly constructed.
@@ 2.2 Spectral stability of Hermitian matrices
Let $\bm{W}$} and $\bm{W}$$^{\prime}$ be Hermitian matrices of size $n$ (i.e. symmetric for real-valued matrices), and let $\delta{\bm{W}}={\bm{W}}^{\prime}-{\bm{W}}$. Also, $\forall i\in[1,\dots,n]$, let $\sigma_{i}({\bm{W}})$ represent the $i$-th largest eigenvalue of $\bm{W}$}. The ordering $\sigma_{1}({\bm{W}})\geq\cdots\geq\sigma_{n}({\bm{W}})$ can always be established as the eigenvalues of Hermitian matrices are guaranteed to be real.
@@
$\forall i\in[1,\dots,n]$, $\sigma_{i}({\bm{W}})$ is Lipschitz-continuous on the space of Hermitian matrices with operator norm:
$|\sigma_{i}({\bm{W}}^{\prime})-\sigma_{i}({\bm{W}})|\leq\max_{i=1,\dots,n}(|\sigma_{i}(\delta{\bm{W}})|)=\max(|\sigma_{1}(\delta{\bm{W}})|,|\sigma_{n}(\delta{\bm{W}})|).$ (7)
@@ 3 The Proposed In-Training Reduction Scheme
Our general pipeline (illustrated in Figure 1) is designed to be applicable to all types of SSMs as it surgically acts on the dynamical systems within SSM layers of models regardless of the choice of projections, non-linear activations, convolutions, skip connections, etc.
@@ 3.1 CompreSSM: the algorithm
At a given training step, we proceed per block. For an input feature vector sequence ${\bm{x}}\in\mathbb{R}^{H\times L}$, where $H$ is the inner dimension and $L$ the sequence length, common SSM blocks contain either a single Multi-Input Multi-Output (MIMO) system that transforms the sequence ${\bm{x}}\in\mathbb{R}^{H\times L}$ to ${\bm{y}}\in\mathbb{R}^{H\times L}$, or $H$ independent per-channel Single-Input Single-Output (SISO) systems that transform ${\bm{x}}^{i}\in\mathbb{R}^{1\times L}$ to ${\bm{y}}^{i}\in\mathbb{R}^{1\times L}$, for $i\in\{1,\dots,H\}$. In the latter case, we proceed per channel.
@@
    Find the smallest rank $r$ such that the top-$r$ singular values retain a predetermined threshold $(1-\tau)\in[0,1]$ of the total energy,
    $r=\min_{k}\Bigl\{\sum_{i=1}^{k}\sigma_{i}\;\geq\;(1-\tau)\sum_{i=1}^{n}\sigma_{i}\Bigr\},$ (8)
@@ 3.3 In-Training Reduction
Also recall that the Hankel singular values can be obtained as the square root of the eigenvalues of ${\bm{P}}{\bm{Q}}$ (Equation 4). Generally, ${\bm{P}}{\bm{Q}}$ need not be symmetric, but we use this form for computational efficiency. Noticing that ${\bm{P}}{\bm{Q}}$ is similar to the symmetric positive definite matrix ${\bm{P}}^{\frac{1}{2}}{\bm{Q}}{\bm{P}}^{\frac{1}{2}}$, we let ${\bm{H}}=({\bm{P}}^{\frac{1}{2}}{\bm{Q}}{\bm{P}}^{\frac{1}{2}})^{\frac{1}{2}}$. The matrix $\bm{H}$}’s eigenvalues are exactly the Hankel singular values. In addition, by composition, $\bm{H}$} is continuous with respect to the perturbations to the system, so we write ${\bm{H}}^{\prime}={\bm{H}}+\delta{\bm{H}}$ with $\delta{\bm{H}}$ a continuous function of $(\delta{\bm{A}},\delta{\bm{B}},\delta{\bm{C}})$.
@@ 4.1 Experimental Setup
Reduction starts from the full model order reported in column 4 of Table 3. For all datasets but IMDB and sMNIST, we attempt four equidistant reduction steps during the warm-up period of the learning rate, which equals $10\%$ of the total steps. Doing so ensures maximal speed up potential for the subsequent $90\%$ of training, while also staying robust to large early training perturbations.
@@ 4.2 Empirical Results
However, on datasets where state dimension does correlate with model performance, CompreSSM improves small model performance. On CIFAR10, for example, model performance almost stays constant as a function of reduction tolerance (and thus for different state dimensions), while the non-compressed counterparts exhibit an approximately $10\%$ performance drop (see entry 1 in Table 1 or Figure 3(a)). The sMNIST results paint a similar picture (entry 5 in Table 1 or Figure 4(a)).
@@
On Pathfinder, we can observe a similar trend (Figure 4(e)), where the unreduced model performance does not show a strong correlation with the state dimension. Just for the smallest state dimension, the unreduced model sees a small drop in performance and is outperformed by the top-1 reduced model.
The IMDB results reveal the importance of the prerequisites mentioned in Section 3.1 (see Figure 4(d)). Even after significantly reducing the parameters and increasing the droprate (see Appendix A), the unreduced models only learn for roughly $8k$ steps before overfitting. However, in order to do in-training balanced truncation, the actual training phase needs to be long enough to allow for a couple of reduction steps with a large enough spacing so that the model can follow the training dynamics for a while without being pruned. Indeed, for non-aggressive pruning (that is, pruning with small tolerance $\tau$), the top reduced models often outperform the baseline.
@@ 5 Related Work
Though not a compression strategy, the authors in Yu et al. (2024) approach the problem of HSV efficiency from the perspective of initialization, aiming to avoid fast decaying singular values.
@@ 6 Conclusion
We propose CompreSSM: a framework for principled in-training compression of SSMs. Leveraging control theory, we utilize Weyl’s theorem to track the evolution of HSVs during learning. Empirically, we observe that the relative contributions of individual state components remain highly stable. This allows us to safely apply balanced truncation early in the training process.
CompreSSM is put to the test with LRUs on a range of LRA tasks. Provided there is a correlation between state capacity and model performance, we verify empirically that compressed models outperform their uncompressed counterparts while delivering better performance per unit of training time. In short, CompreSSM enables performance of high capacity models at both the inference and training overheads of smaller counterparts.
Finally, we provide an implementation strategy for linear time-varying systems. In the selective case, we propose averaging the dynamics over the input space before applying the reduction scheme. Future work will focus on extending CompreSSM to linear attention models, e.g., Gated Linear Attention (Yang et al., 2023), Mamba2 (Dao and Gu, 2024), and Gated DeltaNet (Yang et al., 2024).
Acknowledgments
Philipp Nazari was supported by the Max Planck ETH Center for Learning Systems. This work was supported in part by the Hector Foundation, the Boeing Company, and the ONR Science of Autonomy program N00014-23-1-2354.
@@ C.1 Balanced Truncation Sanity Check
Figure 5 shows the average model accuracy on the validation dataset during training, and Figure 6 depicts the HSVs evolution for a single seed across the three reduction schemes. Removing the top HSVs clearly incurs catastrophic performance damage, that is not recoverable even with 90% of subsequent training. These drops in performance are present yet less pronounced when removing random HSVs. In this case, the original gains gathered from training at larger state dimension are not completely lost, as the network is somewhat able to recover to perform slightly better than the baseline trained with dimension 32 from the start. In the case of proper bottom HSV truncation, there are no notable drops in performance in comparison to the 256 baseline, apart from a small departure at the final reduction step. The performance observations are also backed up by the HSVs evolution in each case. Indeed, we observe that the correct balanced truncation maintains a majority concentration of large HSVs throughout training, all within a single order of magnitude, maximizing the per dimension contribution. Random truncation leads to a sparser spread of HSVs with many small surviving HSVs trailing an order of magnitude or two behind the leading dimensions. Expressivity is most diluted in the adversarial case, where removing the top values leads to dynamics with only a small proportion of HSVs carrying training while the largest bulk remains several orders of magnitude behind.
@@ D.2 Reduction Pipeline Complexity
For ${\bm{B}}\in\mathbb{R}^{n\times p}$, computing ${\bm{B}}{\bm{B}}^{\top}$ takes $\mathcal{O}(n^{2}p)$ FLOPS. Similarly, computing ${\bm{C}}^{\top}{\bm{C}}$ for ${\bm{C}}\in\mathbb{R}^{q\times n}$ takes $\mathcal{O}(n^{2}q)$ FLOPS. Thus, computing the Gramians $\bm{P}$} and $\bm{Q}$} requires $\mathcal{O}(n^{2}p+3n^{2})$ and $\mathcal{O}(n^{2}q+3n^{2})$ FLOPS, respectively.
@@
Step 4: finding the smallest rank $r$ that maintains at least $(1-\tau)$ of the total energy requires computing the cumulative sum of the HSV, which is $\mathcal{O}(n)$ FLOPS.
@@ E.3.2 Practical reduction workflow
From each surrogate we form controllability and observability Gramians ${\bm{P}}^{i}$ and ${\bm{Q}}^{i}$. Their product yields the Hankel singular values ${\bm{\sigma}}^{i}$, which determine the retained dimensionality $r^{i}$. We compute the balancing transform ${\bm{T}}^{i}$ and its inverse $({\bm{T}}^{i})^{-1}$ that simultaneously diagonalize the Gramians ${\bm{P}}^{i}$ and ${\bm{Q}}^{i}$.
Applying $({\bm{T}}^{i})^{-1}\widetilde{{\bm{A}}}^{i}{\bm{T}}^{i}$ and truncating to the first $r^{i}$ rows/columns produces ${\bm{A}}_{\text{red}}^{i}$. Similarly,
${\bm{B}}_{\text{red}}^{i}=({\bm{T}}^{i})^{-1}\widetilde{{\bm{B}}}^{i}\big[:r^{i}\big],\qquad{\bm{C}}_{\text{red}}^{i}=\widetilde{{\bm{C}}}\,{\bm{T}}^{i}\big[:,:r^{i}\big].$
@@
We then keep only the first $r^{i}$ coordinates before feeding the selective recurrence. The vector of ranks ${\bm{r}}=(r^{1},\dots,r^{d_{\text{inner}}})$ is passed to the selective-scan CUDA kernels, which terminate their inner loops at $r^{i}$ for channel $i$. This eliminates all work on pruned coordinates and makes the runtime cost proportional to the reduced dimension.
@@ E.3.3 Experimental Evaluation:
We test the performance of CompreSSM in this setting by compressing with $\tau=0.001$. This means the model retains at least $99.99\%$ of its energy at every reduction step. The final model has an average state dimension of $\sim 12$, starting from $128$. This shows that the HSV spectrum is very tail-heavy, with most energy in just a handful of dimensions. This explains why random dropping of HSV (the red markers in Figure 18) performs competitively with balanced truncation (the orange markers), though at a higher variance.
Even though the correlation between state dimension and performance is not very strong, CompreSSM yields competitive performance on Mamba, at a lower variance, especially when accounting for outliers. However, we believe the approach for LTV systems should be refined in later works to account for heavy-tailed distributions of HSV.